OpenStudy (anonymous):

Write each fraction in simplest form. 4/20

OpenStudy (bibby):

blaze it

OpenStudy (ybarrap):

4/20 = 2*2/2*10 = 2/10 = 2*1/2*5 = 1/5

OpenStudy (anonymous):

Thank you @ybarrap

OpenStudy (ybarrap):

NP

OpenStudy (anonymous):

1/5

OpenStudy (anonymous):

1/5 4 goes into 4 1 time and into 20 5 times

OpenStudy (anonymous):

1\5